# An Empirical Study of Federated Prompt Learning for Vision Language Model

### v1

@@ 3.2.3 Federated Prompt Learning
[…]
@@ 4.1 Datasets
  • Cifar-100 Krizhevsky and Hinton (2009) is a famous classification dataset, containing 32 $\times$ 32 images of 100 classes. Training and validating sets are composed of $50,\!000$ and $10,\!000$ images.
  • DomainNet Peng et al. (2019) contains $6$ domains: Sketch (S), real (R), painting (P), infograph (I), quickdraw (Q), and clipart (C) with $365$ classes. It consists $48k-172k$ images per domain.
@@ 4.2 Implementation Details
  • Network Structure: We employ the pre-trained CLIP model Radford et al. (2021) with a ViT-B/16 image encoder backbone, using the official implementation from the repository [^1] https://github.com/openai/CLIP. For prompt construction, both for the text and visual encoders, we refer to the open-source implementations of CoOPZhou et al. (2022), Jia et al. (2022), and PromptSRC Khattak et al. (2023b) available at [^2] https://github.com/KaiyangZhou/CoOp, [^3] https://github.com/KMnP/vpt, [^4] https://github.com/muzairkhattak/PromptSRC. The prompt length for both visual and textual prompts is set to $L=16$. Visual prompts $\mathbf{P}_{v}$ are inserted into the first 9 layers of the transformer. Both textual and visual prompts are randomly initialized with the normal distribution.
@@
  • Training Settings: To ensure a fair comparison, we follow Huang et al. (2023b) and set the communication epoch $T=50$ and the local update round $E=1$ for all settings. The training batch size is $32$, and we use SGD as the local update optimizer. The corresponding weight decay is $\eta=1e-5$ and momentum is set to $0.9$. The local client learning rate is $0.001$ for all scenarios. To mimic the label skew, we follow the common experiments setting and use Dirichlet sampling Balakrishnan et al. (2019). We fix the random seed at 00 to ensure reproducibility and run the experiments on NVIDIA 3090.
@@ 5.1 Behavior Discrepancy between LPT and VPT in Label Skew and Domain Shift
Regarding Q1 in Sec. 1, we can obtain the answer: A1: There exists Behavior Discrepancy between LPT and VPT under label skew and domain shift. We restate this key phenomenon as follows: {mdframed}[backgroundcolor=cadetblue!10, linewidth=0.8pt, linecolor=cadetblue!80, roundcorner=5pt] Label Skew Prompt Consistency:
For label skew, VPT exhibits consistent optimization across clients and with the global server, leading to unified visual mapping.
{mdframed}
[backgroundcolor=cadetblue!10, linewidth=0.8pt, linecolor=cadetblue!80, roundcorner=5pt] Domain Shift Prompt Discrepancy:
For domain shift, diverse optimization in LPT enables better adaptation to distinct domains.
@@ 5.2 Robustness of Federated Prompt Learning Under Various Settings
Our experiments on Cifar-100 and Office-Home shown in Fig. 5 reveal an interesting pattern: In label skew scenarios, weighted aggregation leads to better performance, as larger clients contribute more meaningfully to the global prompt optimization. In domain shift scenarios, equal aggregation performs better, as it prevents overfitting to the data-rich clients and helps maintain diversity across domains. However, it is important to note that the overall difference in performance between these aggregation methods is relatively small. More importantly, both methods still outperform CLIP’s zero-shot inference, reinforcing the robustness of federated prompt learning to different aggregation strategies. ggests that while aggregation choice may slightly influence results, FPL remains stable and effective across different settings.

### v2

@@ 3.2.3 Federated Prompt Learning
$\mathbf{P}^{m}\leftarrow{}\mathbf{P}^{g}$ $\text{Distribution},$ (8)
$\text{arg}\min_{\mathbf{P}^{m}}\mathbb{E}_{(x,{y})\sim D^{m}}\,\mathcal{L}(\mathbf{{z}},\mathbf{\tilde{w}})/\mathcal{L}(\mathbf{\tilde{z}},\mathbf{{w}})$ $\text{Optimization},$
$\mathbf{P}^{g}=\sum_{m}^{M}\frac{N^{m}}{N}\mathbf{P}^{m}$ $\text{Aggregation},$
@@ 4.1 Datasets
  • Cifar-100 Krizhevsky and Hinton (2009) is a famous classification dataset, containing 32 $\times$ 32 images of 100 classes. Training and validating sets are composed of 50,00050,\!000 and 10,00010,\!000 images.
  • DomainNet Peng et al. (2019) contains $6$ domains: Sketch (S), real (R), painting (P), infograph (I), quickdraw (Q), and clipart (C) with $365$ classes. It consists 48k172k48k-172k images per domain.
@@ 4.2 Implementation Details
  • Network Structure: We employ the pre-trained CLIP model Radford et al. (2021) with a ViT-B/16 image encoder backbone, using the official implementation from the repository [^1] https://github.com/openai/CLIP. For prompt construction, both for the text and visual encoders, we refer to the open-source implementations of CoOPZhou et al. (2022), VPTJia et al. (2022), and PromptSRC Khattak et al. (2023b) available at [^2] https://github.com/KaiyangZhou/CoOp, [^3] https://github.com/KMnP/vpt, [^4] https://github.com/muzairkhattak/PromptSRC. The prompt length for both visual and textual prompts is set to $L=16$. Visual prompts $\mathbf{P}_{v}$ are inserted into the first 9 layers of the transformer. Both textual and visual prompts are randomly initialized with the normal distribution.
@@
  • Training Settings: To ensure a fair comparison, we follow Huang et al. (2023b) and set the communication epoch $T=50$ and the local update round $E=1$ for all settings. The training batch size is $32$, and we use SGD as the local update optimizer. The corresponding weight decay is $\eta=1e-5$ and momentum is set to $0.9$. The local client learning rate is $0.001$ for all scenarios. To mimic the label skew, we follow the common experiments setting and use Dirichlet sampling Balakrishnan et al. (2019). We fix the random seed at 0 to ensure reproducibility and run the experiments on NVIDIA 3090.
@@ 5.1 Behavior Discrepancy between LPT and VPT in Label Skew and Domain Shift
Regarding Q1 in Sec. 1, we can obtain the answer: A1: There exists Behavior Discrepancy between LPT and VPT under label skew and domain shift. We restate this key phenomenon as follows: Label Skew Prompt Consistency: For label skew, VPT exhibits consistent optimization across clients and with the global server, leading to unified visual mapping.
Domain Shift Prompt Discrepancy: For domain shift, diverse optimization in LPT enables better adaptation to distinct domains.
@@ 5.2 Robustness of Federated Prompt Learning Under Various Settings
Our experiments on Cifar-100 and Office-Home shown in Fig. 5 reveal an interesting pattern: In label skew scenarios, equal aggregation leads to better performance, as treating clients uniformly prevents bias toward label-rich participants and ensures balanced optimization across classes. In domain shift scenarios, weighted aggregation performs better, as emphasizing larger domains helps the global model capture the dominant distribution while still retaining cross-domain knowledge. However, it is important to note that the overall difference in performance between these aggregation methods is relatively small. More importantly, both methods still outperform CLIP’s zero-shot inference, reinforcing the robustness of federated prompt learning to different aggregation strategies. ggests that while aggregation choice may slightly influence results, FPL remains stable and effective across different settings.